OpenStudy (anonymous):

HELP PLEASE! Peanuts cost $1.20 per lb. and cashews cost $4.40 per lb. A ten lb. mixture selling for $2 per pound should conatin how many lbs. of each type of nut?

OpenStudy (amistre64):

p + c = 10 1.20p + 4.40c = ... hmmm

OpenStudy (amistre64):

prolly should account for the 10 lbs p + c = 10 12p + 44c = 20

OpenStudy (anonymous):

lets say that we mixed X lbs of peanuts and Y lbs of Cashews. Then, according to question. our total pound is 10 so, X + Y = 10 ---------(i) again, total price of the 10 pounds mixture is 10 * 2 =20 but the total price by mixing should be, 1.2X + 4.4 Y = 20 --------(ii) Now, solve --(i) and ---(ii) and get the values of X and Y which are your required answer

OpenStudy (amistre64):

i think you ii is off

OpenStudy (amistre64):

so is mine tho

OpenStudy (anonymous):

@amistre64 I think yours one is off.. mine is okay.. just do the dimensional check for verification (like we do in Physics), per pound's price multiplied by the no. of pounds should give the total price.. not the per pound price again...

OpenStudy (amistre64):

i may be seeing it off: cost X + costY = 2 dollars p + c = 10 ; 10 lbs 1.2p + 4.4c = 2 ; 10lbs for 2 dollars

OpenStudy (amistre64):

ahhh, i see it now; 10 lbs for 2/lb

OpenStudy (anonymous):

Joe invested part of $8000 at 11% annual interest. The rest was invested at 9%,last years interest on the $8000 was $850. How much did she invest at each rate?